# Lived Experience Not Found: LLMs Struggle to Align with Experts on Addressing Adverse Drug Reactions from Psychiatric Medication Use

### v1

@@ 1 Introduction
Given the current limitations of healthcare systems and social media platforms, alongside the growing capabilities of LLMs in mental health-related tasks Yang et al. (2023, 2024); Singhal et al. (2023b), LLMs have the potential to bridge the gap in online discussions by providing high-quality, contextual responses to ADR queries related to psychiatric medications. While previous studies have focused on detecting ADRs using deep learning methods, these efforts have primarily addressed non-mental health-related ADRs Mesbah et al. (2019a); Sarker and Gonzalez (2015); Karimi et al. (2015). Detecting ADRs related to psychiatric medications and evaluation of feedback responses from LLMs towards ADR-related queries, remains unexplored. This gap is particularly significant because addressing ADRs caused by psychiatric medications presents unique challenges, given the complex interplay between mental health conditions, their symptoms, and the potential for psychiatric medications to either alleviate or exacerbate those issues. Furthermore, for LLMs to meaningfully contribute to online discussions and effectively address ADR-related queries, it is essential to rigorously evaluate the quality of their long-form text responses and their alignment with expert knowledge in specialized domains such as psychiatry. This evaluation must consider the LLMs’ ability to grasp the complexities of psychiatric medication ADRs and deliver responses that are contextually nuanced. In response to these needs and challenges, we address the following research questions:
@@
RQ1 results show that both ADR detection and ADR multiclass classification are challenging tasks, with the top model in a few-shot setting achieving F1 scores of 75.38 and 76.69 in respective tasks. We observed that all models exhibited a “risk-averse” nature, leading to a false-positive rate of over 70%. Additionally, models struggled with non-dose-related and time-related ADRs, with GPT-4 Turbo misclassifying 51% and 50% of these instances, highlighting difficulties in grasping nuanced ADR types. For RQ2, LLM-generated responses were significantly harder to read than expert-written responses. In contrast, there was no observed significant difference in emotional or tone alignment between LLM and expert responses. However, the best model (OpenBioLLM-70B) achieved only 70.86% alignment with expert harm-reduction strategies, and LLMs provided 12.32% less actionable advice on average. Given the observations, our research has important real-world implications. The proposed benchmark provides a resource for evaluating LLMs on tasks involving the interaction between mental health conditions and psychiatric medications. The proposed framework hold practical utility for policymakers, practitioners, and healthcare professionals to assess LLM performance, especially in strategy-driven tasks in high-risk domains.
@@ 5.1 Zero-shot prompting on Psych-ADR
Models exhibited a “risk-averse” tendency, and prone to commit false-positive errors. In both ADR detection and multiclass classification tasks, all models displayed “risk-averse" behavior, often mislabeling posts without ADRs as positive for ADRs. In zero-shot settings, Claude 3 Opus had a false-positive rate of 42% for ‘ADR-No’ labels, while Claude 3 Haiku’s false positive rate was as high as 97% (see Appendix F). Similarly, in ADR multiclass classification, models struggled to distinguish between non-dose-related, dose-related, and time-related ADRs. GPT-4 Turbo misclassified 51% of non-dose-related and 50% of time-related ADRs in zero-shot settings. This risk-averse tendency indicates a lack of nuanced understanding of ADR complexities, which could lead to (a) patients discontinuing treatment Horne and Weinman (1999); Horne et al. (2005), (b) increased fear about their conditions Starcevic and Berle (2013), and (c) “alert-fatigue” among healthcare providers Phansalkar et al. (2013).
@@ 6 RQ2: Alignment between human and AI feedback
We then evaluated alignment of HRS for each LLM-expert response pair using two methods. First, we used AlignScore Zha et al. (2023), a widely-used text alignment method providing a score between 0 and 1 based using a fine-tuned RoBERTa-large model Liu et al. (2019). We computed AlignScore for each strategy from the LLM response against the expert response. We obtained a response-level AlignScore by averaging the scores across all HRS for the response. Second, for a more interpretable alignment score, we prompted GPT-4o with in-context examples to reason and classify if a strategy is aligned with the expert’s response. We computed a response-level GPT-4o score by computing the percentage of aligned HRS over total number of HRS. These two approaches ensured robustness by covering both a continuous alignment score and one based on reasoned binary alignment labels. We conducted another round of human evaluation for the GPT-4o score, where four annotators annotated 40 responses, achieving a 95% correlation with GPT-4o’s score and reasoning. Prompts for LLM-based tasks are presented in Table 12, 13 & 14, and human evaluation details are presented in Appendix I.1.
@@ 6.1 Results
Emotional and tone alignment. Figure 2 presents the mean KL-divergence score for the distribution of 11 Empath categories between LLM responses and expert-written response. A $\chi^{2}$ test was conducted to assess the differences in category distributions, and the $p$-values were non-significant across all models, indicating that the models’ responses were not significantly different from the expert-written responses in terms of emotions expressed and the tone used. Further, we observed that larger and more capable models from the Llama and Claude families showed greater alignment with expert responses across different emotional and tone related categories. Interestingly, Llama-3 Med42v2 70B performed the worst. This could be attributed to the fact that a major portion of dataset used for instruction fine-tuning for this model was obtained from the medical and biomedical literature, which may not prioritize emotional communication while providing responses Christophe et al. (2024).
@@ Appendix D Annotation Task Details
Figure 5 presents the sample of structure of the expert responses provided in the Psych-ADR benchmark. The most experienced doctor on the collaborating team provided the responses, with each taking $\sim$8 minutes to answer on average.

### v2

@@ 1 Introduction
Given the current limitations of healthcare systems and social media platforms, alongside the growing capabilities of LLMs in mental health-related tasks Yang et al. (2023, 2024); Singhal et al. (2023b), LLMs have the potential to bridge the gap in online discussions by providing high-quality, contextual responses to ADR queries related to psychiatric medications. While previous studies have focused on detecting ADRs using deep learning methods, these efforts have primarily addressed non-mental health-related ADRs Mesbah et al. (2019a); Sarker and Gonzalez (2015); Karimi et al. (2015). Detecting ADRs related to psychiatric medications and evaluation of feedback responses from LLMs towards ADR-related queries, remains unexplored. This gap is particularly significant because addressing ADRs caused by psychiatric medications presents unique challenges, given the complex interplay between mental health conditions, their symptoms, and the potential for psychiatric medications to either alleviate or exacerbate those issues. Furthermore, for LLMs to meaningfully contribute to online discussions and effectively address ADR-related queries, it is essential to rigorously evaluate the quality of their long-form text responses and their alignment with expert knowledge in specialized domains such as psychiatry. This evaluation must consider the LLMs’ ability to grasp the complexities of psychiatric medication ADRs and deliver responses that are contextually nuanced. Additionally, there is a need to evaluate the ability of LLMs to accurately portray the lived experiences [^2] Personal knowledge about the world gained through direct, first-hand involvement in everyday events rather than through representations constructed by other people. of healthcare providers in addressing ADR-related queries which has been considered as an important factor in providing effective mental health support. In response to these needs and challenges, we address the following research questions:
@@
RQ1 results show that both ADR detection and ADR multiclass classification are challenging tasks, with the top model in a few-shot setting achieving F1 scores of 75.38 and 76.69 in respective tasks. We observed that all models exhibited a “risk-averse” nature, leading to a false-positive rate of over 70%. Additionally, models struggled with non-dose-related and time-related ADRs, with GPT-4 Turbo misclassifying 51% and 50% of these instances, highlighting difficulties in grasping nuanced ADR types. For RQ2, LLM-generated responses were significantly harder to read than expert-written responses. In contrast, there was no observed significant difference in emotional or tone alignment between LLM and expert responses. However, the best model (OpenBioLLM-70B) achieved only 70.86% alignment with expert harm-reduction strategies, and LLMs provided 12.32% less actionable advice on average. Given the observations, our research has important real-world implications. The proposed benchmark provides a resource for evaluating LLMs on tasks involving the interaction between mental health conditions and psychiatric medications. The proposed framework hold practical utility for policymakers, practitioners, and healthcare professionals to assess LLM performance, especially in strategy-driven tasks in high-risk domains. The code repository for this work can be found here [^3] https://github.com/mohit3011/Lived-Experience-Not-Found.
@@ 5.1 Zero-shot prompting on Psych-ADR
Models exhibited a “risk-averse” tendency, and prone to commit false-positive errors. In both ADR detection and multiclass classification tasks, all models displayed “risk-averse" behavior, often mislabeling posts without ADRs as positive for ADRs (see Appendix G for qualitative analysis). In zero-shot settings, Claude 3 Opus had a false-positive rate of 42% for ‘ADR-No’ labels, while Claude 3 Haiku’s false positive rate was as high as 97% (see Appendix F). Similarly, in ADR multiclass classification, models struggled to distinguish between non-dose-related, dose-related, and time-related ADRs. GPT-4 Turbo misclassified 51% of non-dose-related and 50% of time-related ADRs in zero-shot settings. This risk-averse tendency indicates a lack of nuanced understanding of ADR complexities, which could lead to (a) patients discontinuing treatment Horne and Weinman (1999); Horne et al. (2005), (b) increased fear about their conditions Starcevic and Berle (2013), and (c) “alert-fatigue” among healthcare providers Phansalkar et al. (2013).
@@ 6 RQ2: Alignment between human and AI feedback
We then evaluated alignment of HRS for each LLM-expert response pair using two methods. First, we used AlignScore Zha et al. (2023), a widely-used text alignment method providing a score between 0 and 1 based using a fine-tuned RoBERTa-large model Liu et al. (2019). We computed AlignScore for each strategy from the LLM response against the expert response. We obtained a response-level AlignScore by averaging the scores across all HRS for the response. Second, for a more interpretable alignment score, we prompted GPT-4o with in-context examples to reason and classify if a strategy is aligned with the expert’s response. We computed a response-level GPT-4o score by computing the percentage of aligned HRS over total number of HRS. These two approaches ensured robustness by covering both a continuous alignment score and one based on reasoned binary alignment labels. We conducted another round of human evaluation for the GPT-4o score, where four annotators annotated 40 responses, achieving a 95% correlation with GPT-4o’s score and reasoning. Prompts for LLM-based tasks are presented in Table 13, 14 & 15, and human evaluation details are presented in Appendix I.1.
@@ 6.1 Results
Emotional and tone alignment. Figure 2 presents the mean KL-divergence score for the distribution of 8 Empath categories between LLM responses and expert-written response. A $\chi^{2}$ test was conducted to assess the differences in category distributions, and the $p$-values were non-significant across all models, indicating that the models’ responses were not significantly different from the expert-written responses in terms of emotions expressed and the tone used. Further, we observed that larger and more capable models from the Llama and Claude families showed greater alignment with expert responses across different emotional and tone related categories. Interestingly, Llama-3 Med42v2 70B performed the worst. This could be attributed to the fact that a major portion of dataset used for instruction fine-tuning for this model was obtained from the medical and biomedical literature, which may not prioritize emotional communication while providing responses Christophe et al. (2024).
@@ 7 Related Work
Importance of lived experience: The importance of lived experience has been studied in various works across the field of mental health research, psychology, and education. Understanding lived experiences provides insight into individuals’ personal realities and preferences, contributing towards a deeper understanding of their experiences, expectations and requirements. In mental health research, previous studies have highlighted the importance of understanding the experiences and realities of individuals living with mental health conditions for providing better treatment Gilbert and Stickley (2012); Repper and Carter (2011)Byrne et al. (2013) further highlighted that students showed positive attitudes and increased self-awareness towards the impact of mental illness on individuals when they were taught by people with lived experience of mental health challenges. Past research in psychology has also stressed on the understanding of the lived experience of individuals belonging to different backgrounds. Previous studies have also highlighted the importance of lived experience in the form of experiential knowledge among the healthcare provider for making decisions Lyu et al. (2023); Palukka et al. (2021). In summary, previous research in mental health and psychology has emphasized on the multifaceted importance of lived experience among the patients, educators and healthcare providers.
@@ 9 Broader Implications
Disentangling inclusion of humanistic features in LLMs and advocacy for inclusion of lived experience. While our work provides evidence of the lack of lived experience, which is essential for understanding the nuances of a complex task such as ADR detection and for proposing mitigation strategies in model responses, we do not advocate for increasing human-like features in LLMs. Previous studies have suggested that heightened anthropomorphism, independent of whether it is accompanied by enhanced capabilities, can increase trust among individuals Natarajan and Gombolay (2020); Chen and Park (2021). Hence, developers and researchers need to be cautious before introducing such features as individuals may trust LLM responses even when they provide incorrect or inconsistent information which can be hazardous in high-risk domains such as healthcare. In contrast, we advocate for approaches that align with previous research, which has shown that the efficacy of LLMs in the healthcare domain can be enhanced through fine-tuning on specialized data or by incorporating useful features into the model, without introducing human-like features Belyaeva et al. (2023); Li et al. (2023).
10 Limitations
@@ 11 Ethical Considerations
12 Acknowledgment
Chandra and De Choudhury were partly supported through National Science Foundation (NSF) grant #2230692, a grant from the American Foundation for Suicide Prevention (AFSP), and Azure compute resources from Microsoft. The findings, interpretations, and conclusions of this paper are those of the authors and do not represent the official views of NSF, AFSP or Microsoft. We further acknowledge the contributions from Dr. Santiago Alvarez Lesmes towards data annotation and assessment.
@@ Appendix D Annotation Task Details
D.1 ADR Post Reply Annotation and Generation
Figure 5 presents the sample of structure of the expert responses provided in the Psych-ADR benchmark. The most experienced doctor on the collaborating team provided the responses, with each taking $\sim$8 minutes to answer on average. We provided the same set of instructions to the domain expert and the LLMs for writing the response to an ADR related query. Both the experts and the LLMs were prompted to write responses that follow a strict logical structure typically seen in clinical settings:
  • Start with empathizing with the post writer and acknowledging their concern.
  • As step 2, provide a brief explanation for the potential diagnosis as an adverse drug reaction.
  • As step 3, talk briefly about the evidence of the adverse drug reaction which is available in the post and provide a brief reasoning for proposing a harm reduction strategy.
  • As step 4, ask potential questions to get additional information that might help in providing a better response.
  • As step 5, provide a brief explanation of the harm reduction strategy that the poster should follow to get relief.
  • Finally, ask the post writer about how they feel about the proposed harm reduction strategy.
Table 10 presents the prompt for ADR reply prompt generation.
Type Prompt
System Prompt You are an expert psychiatrist and psychologist, and you help people with their mental health and adverse drug reaction (ADR) related queries. You provide advice, and guidance to people who are experiencing mental health issues. Given below is the guideline for generating an ideal reply (Presented as IDEAL_REPLY_TEMPLATE), the reddit post title (Presented as POST_TITLE), and the reddit post text (Presented as POST_TEXT) expressing adverse drug reaction (ADR) related to a psychiatric medication/s.
IDEAL_REPLY_TEMPLATE: An "ideal reply" follows the below steps:
1. Start with empathizing with the post writer and acknowledging their concern.
2. As step 2, provide a brief explanation for the potential diagnosis as an adverse drug reaction.
3. As step 3, talk briefly about the evidence of the adverse drug reaction which is available in the post and provide a brief reasoning for proposing a harm reduction strategy.
4. As step 4, ask potential questions to get additional information that might help in providing a better response.
5. As step 5, provide a brief explanation of the harm reduction strategy that the poster should follow to get relief.
6. Finally, ask the post writer about how they feel about the proposed harm reduction strategy.
Below are some additional guidelines:
1. Do not words like dear poster, dear user, best regards, etc. in the response.
2. Some of the steps of an "ideal answer" are optional and you can skip those if that makes more sense, but it is recommended to include them.
3. The response should be concise and to the point and with the word limit of 225 words or 300 token. Never exceed this word/token limit.
Keeping the context of the POST_TITLE and POST_TEXT in view and using the guideline provided in IDEAL_REPLY_TEMPLATE, your task is to generate a reply to the post. Your response should be helpful and aim to provide a solution to the issues/problems mentioned in the post.
User Prompt POST_TITLE: <post_title>
POST_TEXT: <post_text>
Table 10: Prompt used for generating the reply for post expressing ADR related concerns.